Electronic Journal: Southwest Journal of Pure and Applied Mathematics Internet: http://rattler.cameron.edu/swjpam.html ISSN 1083-0464 Vol. No. 01, 1997, pp. 13–17. Submitted: January 1, 1997. Published: June 29, 1997.

ESTIMATION OF CHANNEL HOLDING TIME

DISTRIBUTION IN CELLULAR TELEPHONE SYSTEMS

M.H. Cho^{1}, K.R. Cho^{2}, K.S. Choi^{2}, F. Szidarovszky^{3}, Y. You^{1}

Abstract. In estimating the channel holding time distribution in cellular telephone systems, the pdf of the fraction of two non-negative random variables, distance and speed, has to be determined. Instead of the repeated application of numerical integration we show that the pdf of the holding time is the solution of an initial value problem of a non-linear ordinary differential equation, which can be solved very efficiently.

1991 A.M.S. (MOS) Subject classification Codes. 94A40.

Key Words and Phrases. Cellular telephone system, holding time distribution, transformed random variable.

2. Mobile Telecommunication Technology Division, Electronics and Telecommunications Research Institute, Yusong P.O. Box 106, Taejon, 305-600, Korea, E-mail address: mhcho@etri.re.kr

1. Department of Computer and Communication Engineering, Chungbuk National University, Gaesin San 48, Cheongju, Chungbuk, 360-763, Korea

3. Systems& Industrial Engineering, University of Arizona, Tucson, Arizona 85721, U.S.A. E-mail address: szidar@sie.arizona.edu \begin{matrix}c \\ ◯ \end{matrix}1997 Cameron University

Typeset by AMS-TEX